40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Mar 4 10:54:12 2021, Paco, Anchal, Summary, LSC, POY11 measurement, tried to lock Green Yend laser 20210304_POY11_Spectrum_YARMLocked.pdf20210304_POY11_Spectrum_YARMLocked.tar.gz
    Reply  Thu Mar 4 11:59:25 2021, Paco, Anchal, Summary, LSC, Watchdog tripped, Optics damped back 
       Reply  Thu Mar 4 15:48:26 2021, Koji, Summary, PEM, Watchdog tripped, Optics damped back 
Message ID: 15862     Entry time: Thu Mar 4 11:59:25 2021     In reply to: 15861     Reply to this: 15863
Author: Paco, Anchal 
Type: Summary 
Category: LSC 
Subject: Watchdog tripped, Optics damped back 

Gautam came in and noted that the optics damping watchdogs had been tripped by a >5 magnitude earthquake somewhere off the coast of Australia. So, under guided assistance, we manually damped the optics using following:

  • Using the scripts/SUS/reEnableWatchdogs.py script we re-enabled all the watchdogs.
  • Everything except SRM was restored to stable state.
  • Then we clicked on SRM in SUS-> Watchdogs, disabled the Oplevs, shutdown the watchdog.
  • We changed the threshold for watchdog temporarily to 1000 to allow damping.
  • We enabled all the coil outputs  manually. Then enabled watchdog by clicking on Normal.
  • Once the SRM was damped, we shutdown the watchdog, brought back the threshold to 215 and restarted it.

Gautum also noticed that MC autolocker got turned OFF by me (Anchal), we turned it back on and MC engaged the lock again. All good, no harm done.

ELOG V3.1.3-